OpenStudy (anonymous):

Solve: 1/e^x = 7 A. ln ^1/7 B. ln ^7 C. ln 1/7 D. ln 7

OpenStudy (anonymous):

@satellite73 Can you help?

OpenStudy (anonymous):

7*e^x=1 1/7=e^x do you know what to do next?

OpenStudy (anonymous):

No..

OpenStudy (anonymous):

Take ln of both sides... ln(1/7)=x

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

But is it ln^1/7 or just ln 1/7??

OpenStudy (anonymous):

ln 1/7

OpenStudy (anonymous):

Okay.